OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 3x - 2y = 9 and -3x + 2y = -9

OpenStudy (anonymous):

3x=9+2y x=9+2y/2 putting the value of x in eq.2 -18-6y+4y=-18 -2y=36 y=-18

OpenStudy (anonymous):

same way, u can find value of x,, and btw you must have posted this question in mathematics :)

OpenStudy (anonymous):

@d.t. ,, did u get that 1??

OpenStudy (anonymous):

yeah ofcourse...its the same thing i'm studying nowadays

OpenStudy (anonymous):

no,, that endothermic 1??

OpenStudy (anonymous):

yes thanku

OpenStudy (anonymous):

hey @hkaur this is not elimination method .it is substitution method with which you have solved the question and it is not correct either the correct soln is-[x=9+2y/3] puttng the value in eqn 2nd we get -3*(9+2y)/3+2y=-9 -9-2y+2y=-9 -9=-9 which is a false statement hence there is no soln.... and through alimination method we will subtract the two eqns 3x-2y=9 -3x+2y=-9 ------------ o=0 so no soln is there which is a false statement